# Humanoid Locomotion and Manipulation: Current Progress and Challenges in Control, Planning, and Learning thanks: *co-corresponding authors

### v1

@@ VII-E Representation and Composition for Versatile Skills
A good representation of skills makes it easy to compose tasks. In general, a skill can be represented explicitly as a state-action trajectory that accomplishes a task or implicitly as a network structure and its learned weights. In this subsection, we explore implicit representations for skill composition to achieve versatile, multi-skill tasks.
VII-E1 Skill Composition via Mixture of Experts
@@
Recent learning-based approaches enable smooth transitions between multiple skills. Among them, Mixture of Experts (MOE) is widely used. MOE employs a hierarchical architecture; it first trains multiple distinct skills, implicitly encoded in low-level expert policies, and then learns a high-level policy to select [335] or blend [336] these expert networks. This architecture allows for smooth transitions between skills and facilitates the completion of diverse tasks. However, MOE encounters expert imbalance issues favoring certain experts while degrading others; which could diminish the diversity provided by the experts. Instead of obtaining and then blending multiple policies, structured representations improve memory efficiency and allow a single policy to achieve multiple tasks. Next, we introduce three well-received structured representations: motion representation, goal representation, and state transition representation, all shown in Fig. 15.
VII-E2 Motion Representation
@@
Motion representation extracts the essential features and temporal dependencies of high-dimensional long-horizon motions [337]. Specifically, motion representation encodes high-dimensional motions in a low-dimensional latent space. Such latent-space representations are commonly learned in an unsupervised manner using generative models such as Variational Autoencoders (VAEs) [338] and Generative Adversarial Networks (GANs) [339]. Compared to VAEs, GANs have greater potential to generate realistic motions following the reference data distribution, but they are often difficult to train. The result of learning motion representation is often a model that can generate versatile motions given latent codes. Therefore, the generative model can be easily reused to achieve new downstream tasks by pairing it with a high-level task-specific policy. For example, several studies [340, 341, 342] learn a high-level RL policy to efficiently use a reduced-dimensional latent space motion representation and allow simulated humanoid robots to follow a set of user-commanded tasks.
VII-E3 Goal Representation
@@
Besides motion representation, another approach to learning a single policy for multiple tasks is through the representation of goals. The goal is typically represented as a feature vector, which can be encoded from an image of the scene in its final state, a natural language instruction, or a desired state from observing human demonstrations. This goal representation is often paired with Goal-Conditioned Policies (GCPs) [343]. Unlike standard RL policies that achieve only one task, GCPs achieve multiple tasks within a single general policy conditioned on different goals. GCPs have achieved versatile humanoid skills using IL such as diffusion-model based BC [276] and RL with imitation objectives [313, 271, 344].
VII-E4 State Transition Representation
@@
The latent space can also represent the transition dynamics of a Markov Decision Process (MDP). In this representation, data collected from the MDP are used to train a dynamics model that predicts the transition probability between abstract, compressed representations of MDP states. This learned dynamics model is often referred to as a world model [345]. Sampling from a world model yields imaginary data and can be achieved efficiently in massive parallel. By leveraging the imaginary data, Model-Based RL (MBRL) achieves greater efficiency compared to a typical model-free setting, where interaction data must be obtained from a simulated or real environment. MBRL has shown success in agile motor skills on humanoids [346, 347]. Another approach, TD-MPC2 [348], uses the world model in an MPC fashion, planning actions that lead to imaginary trajectories with high scores. Beyond data efficiency, the world model can mitigate the sim-to-real gap by fine-tuning a small batch of real-world data [349].
@@
Conclusion: Enabling robots to accomplish versatile skills and multiple tasks is one of the main trends in robot skill learning. Whereas obtaining and blending single-skill policies is widely explored, more recent methods put efforts into achieving multiple tasks in a single policy. This requires a more structured representation of the skill motion, the task goal, and/or the environment dynamics. Latent space models, goal-conditioned policies, and world models are promising approaches toward this objective. However, many of these methods are still limited to the computer graphics community and yet to be implemented on humanoid hardware.
@@ VII-F Learning for Humanoid Loco-manipulation
Loco-manipulation skills are challenging for learning-based methods as they often struggle with achieving physically stable contact or precise contact forces.
@@
While many learning-based approaches, such as CooHOI [350], demonstrate humanoid loco-manipulation skills in simulation, the physical interactions with external environments or objects are often oversimplified. As a result, only a few studies demonstrate sim-to-real transfer for loco-manipulation skills. This subsection overviews a comprehensive list of studies on learning humanoid loco-manipulation with hardware implementations. A large part of these studies rely on RL, with a few examples shown in Fig. 16. However, RL for loco-manipulation typically involves complicated reward designs that are fine-tuned for specific environments and tasks [351, 352], see Sec. VII-A. To enable loco-manipulation tasks, these methods define the contact sequence either implicitly using reference trajectory [351, 333] or explicitly via reward design [274], thus increasing the success rate of sim-to-real transfer. To address the uncertainty of the mass and other properties of the manipulated object, most RL approaches rely on domain randomization to provide robustness to object parameters [39].
@@
To achieve various loco-manipulation tasks such as playing soccer that involves ball kicking and falling recovery using hands [275], employing a hierarchical RL structure that manages distinctly trained skills is a viable strategy. However, hierarchical RL has scalability issues, see Sec. VII-E. On the other hand, IL methods, particularly those employing RL with motion imitation, have enabled humanoid loco-manipulation via teleoperation, as outlined in Sec. VII-B. Although teleoperation is not autonomous, it is a crucial intermediate step to collect humanoid data. Such humanoid data has led to promising developments in autonomous loco-manipulation skills [271, 272], with the potential for further expansion to more diverse loco-manipulation tasks.
@@
Conclusion: Although learning-based methods for humanoid loco-manipulation are less developed than model-based methods, their significance should not be overlooked. Learning-based methods are potentially more robust, as they can adapt to unstructured scenarios that model-based methods struggle to explicitly address, such as recovering from a fall in an arbitrary configuration [353]. Moreover, learning-based methods can find emergent behaviors that are challenging for model-based methods [282]. The recent success of quadrupedal loco-manipulation [77, 72, 76] shows significant promise for humanoids. However, transferring these algorithms from quadruped to humanoid is challenging because of the complex dynamics of humanoids, which require enhanced safety measures and precise balance control.
@@ VIII Foundation Models for Humanoid Robots
Foundation Models (FMs) are large pre-trained models using Internet-scale data [354]. Recent progress in FMs such as Large Language Models (LLMs) and Vision-Language Models (VLMs) has demonstrated groundbreaking capabilities in solving a wide range of downstream tasks (through in-context learning or fine-tuning), such as code generation, visual question answer, and video understanding [355]. The common sense reasoning capabilities of FMs have inspired many explorations of their applications in robotics [9, 10]. Despite this growing interest, research on FMs specifically for humanoid robots remains sparse. In this section, we first provide an overview of FMs in the context of general robotics (e.g., mobile manipulation, instead of humanoids) and then explore their potential applications to humanoid robots.
Strategies to leverage FMs for robotics can generally be categorized into two, as shown in Fig. 17. The first strategy (Sec. VIII-A) elicits actionable knowledge from pre-trained LLMs/VLMs for robotic tasks, without additional model fine-tuning. The second strategy (Sec. VIII-B) collects abundant robotic data to fine-tune or co-train a Robot Foundation Model (RFM) that generalizes to control tasks with common sense reasoning capability [356, 357, 358, 359].
@@ VIII-A Applying LLMs/VLMs to Humanoid Robots
Applying LLMs/VLMs to humanoids is still a nascent field. Many works have shown successful deployment of LLMs/VLMs on various robot embodiments such as dexterous hands [360], manipulators [361], mobile manipulators [362], quadrupedal robots [363], and bipedal robots [364].
@@
The task planning capability of FMs has advanced the complexity of tasks that a humanoid can accomplish. For instance, Figure AI demonstrates fast, dexterous manipulation skills selected by a VLM that interprets natural language commands and the surrounding environment. In [366] and [367], pre-trained LLMs are used to select skills and task goals for animated humanoid characters. OmniH2O [271] employs GPT-4 [355] to select autonomous skills such as greeting a human. HYPERmotion [368] applies an LLM to construct task graphs that enable a hybrid wheeled-leg robot to execute complex loco-manipulation tasks. However, using FMs to plan with a fixed skill set is limited in skill versatility. In addition, for complex behaviors such as those in humanoid loco-manipulation tasks, it is essential to allow FMs to author the detailed motions of low-level skills, instead of selecting from existing, relatively-abstract low-level skills.
Thus, many research efforts have focused on identifying the best bridge between FMs and low-level robot skills. For example, researchers have proposed to generate code [361, 369] and reward functions [370, 363, 364] as intermediate representations for bipedal and quadrupedal robots. Compared to selecting existing skills, these intermediate representations provide additional flexibility in adjusting the generated motion. Furthermore, FMs can generate whole-body poses [371, 372] and whole-body contacts [373, 374] for humanoid robots. These techniques allow users to intuitively direct a robot’s behavior through expressive inputs such as natural language, images, or even gestures.
@@ VIII-B Building Humanoid Foundation Models
While most FMs are developed in the vision or language domain, building FMs in the robotics domain for embodied intelligence is a natural extension. Similar to LLM and VLM, robot foundation models (RFMs) are large models trained from Internet-scale robotic datasets. The RFMs often process multimodal inputs (e.g., egocentric images, and natural language as task description) and directly interact with the physical world through robot actions. Leveraging the Internet-scale multimodal dataset, RFMs hold the promise of generalization across diverse tasks and provide a natural interface for human-robot interaction, both essential for real-world robot applications.
However, building large RFMs presents significant challenges. Successful implementations are limited to robots that have stable dynamics and a large amount of high-quality data collected with significant resources [357, 375, 376, 359]. Recent works have extended such RFMs to floating-base robots such as drones and quadruped robots [377]. However, building an RFM for humanoid loco-manipulation remains a challenging endeavor: the inherent instability in locomotion and high-dimensional action space in dexterous hands makes it extremely difficult to collect high-quality data efficiently. In addition, humanoid robots have critical safety constraints that often limit the exploration of their full behavioral capabilities, resulting in a lack of humanoid datasets.
@@
One popular approach to building RFMs is to create a multi-task sensorimotor control policy using a high-capacity model that can consume large amounts of robot data. The trained policy can perform a wide variety of low-level skills, even for multiple robot embodiments. To enable high-level reasoning/planning, this approach often employs a hierarchical framework combining a pre-trained LLM or VLM with the low-level control policy, initialized from scratch. During training, this setup aligns the semantic knowledge of the LLM and VLM with the physical behavior of the control policy, enabling cross-modal capabilities such as language-to-action. Transformer [378] is a common choice for low-level policy due to its scalability. To enable interactive user commands, the low-level policy is conditioned on language and/or image inputs, often tokenized with pre-trained text or image encoders from LLMs or VLMs. For example, RT-1 [356] and VIMA [358] both train a language-conditioned visuomotor policy with a large amount of manipulation data and have demonstrated the ability to perform a wide range of skills. Despite significant research efforts in RFMs, applying them to humanoid robots remains an unexplored research area with limited existing work. Notable examples include HumanVLA [303] and SuperPADL [304], which train a humanoid-specific policy aligned with the latent space of pre-trained VLMs, enabling skills based on image and language inputs.

### v2

@@ VII-E Representation and Composition for Versatile Skills
A good representation of skills makes it easy to compose tasks. In general, a skill can be represented explicitly as a state-action trajectory that accomplishes a task or implicitly as a network structure and its learned weights. In this subsection, we explore implicit representations for skill composition to achieve versatile and multi-skill tasks.
@@
Recent learning-based approaches enable smooth transitions between multiple skills. Among them, Mixture of Experts (MOE) is widely used. MOE employs a hierarchical architecture; it first trains multiple distinct skills, implicitly encoded in low-level expert policies, and then learns a high-level policy to select [346] or blend [47] these expert networks. This architecture allows for smooth transitions between skills and facilitates the completion of diverse tasks. However, MOE encounters expert imbalance issues favoring certain experts while degrading others, which could diminish the diversity provided by the experts. Instead of obtaining and then blending multiple policies, structured representations improve memory efficiency and allow a single policy to achieve multiple tasks. Next, we introduce three well-received structured representations: motion representation, goal representation, and state transition representation, all shown in Fig. 15.
VII-E1 Motion Representation
@@
Motion representation extracts the essential features and temporal dependencies of high-dimensional long-horizon motions [347]. Specifically, motion representation encodes high-dimensional motions in a low-dimensional latent space. Such latent-space representations are commonly learned in an unsupervised manner using generative models such as Variational Autoencoders (VAEs) [348] and Generative Adversarial Networks (GANs) [349]. Compared to VAEs, GANs have greater potential to generate realistic motions following the reference data distribution, but they are often difficult to train. The result of learning motion representation is often a model that can synthesize versatile motions given latent codes. Therefore, the generative model can be reused for new downstream tasks by pairing it with a high-level task-specific policy. For example, several studies [350, 351, 352] learn a high-level RL policy to efficiently use a reduced-dimensional latent space motion representation and allow simulated humanoid robots to follow a set of user-commanded tasks.
VII-E2 Goal Representation
@@
Another approach to learning a single policy for multiple tasks is through the representation of goals. The goal is typically represented as a feature vector, which can be encoded from an image of the scene in its final state, a natural language instruction, or a desired state from observing human demonstrations. This goal representation is often paired with Goal-Conditioned Policies (GCPs) [353]. Unlike standard RL policies that achieve only one task, GCPs achieve multiple tasks within a single general policy conditioned on different goals. GCPs have demonstrated versatile humanoid skills using IL, such as diffusion-model based BC [283] and RL with imitation objectives [324, 267, 354].
VII-E3 State Transition Representation
@@
The latent space can also represent the transition dynamics of a Markov Decision Process (MDP). In this representation, data collected from the MDP are used to train a dynamics model that predicts the transition probability between abstract, compressed representations of MDP states. This learned dynamics model is often referred to as a world model [355]. Sampling from a world model yields imaginary data, which can be achieved efficiently in parallel. By leveraging the imaginary data, Model-Based RL (MBRL) achieves greater efficiency compared to a typical model-free setting, where interaction data must be obtained from a simulated or real environment. MBRL has shown success in agile motor skills on humanoids [356, 357]. Another approach, TD-MPC2 [358], uses the world model in an MPC fashion, planning actions that lead to imaginary trajectories with high scores. Beyond data efficiency, the world model can mitigate the sim-to-real gap by fine-tuning a small batch of real-world data [359].
@@
Key Takeaways: Enabling robots to accomplish versatile skills and multiple tasks is one of the main trends in robot skill learning. Whereas obtaining and blending single-skill policies is widely explored, more recent methods put efforts into achieving multiple tasks in a single policy. This requires a more structured representation of the skill motion, the task goal, and/or the environment dynamics. Latent space models, goal-conditioned policies, and world models are promising approaches toward this objective. However, many of these methods are still limited to the computer graphics community and yet to be implemented on humanoid hardware.
@@ VII-F Learning for Humanoid Loco-manipulation
Loco-manipulation skills are challenging for learning-based methods as they often struggle with achieving physically stable contact or precise contact forces. While many learning-based approaches, such as CooHOI [360], demonstrate humanoid loco-manipulation skills in simulation, the physical interactions with external environments or objects are often oversimplified. As a result, only a few studies have demonstrated sim-to-real transfer for loco-manipulation skills. A large part of these studies rely on sim-to-real RL, with a few examples shown in Fig. 16. However, sim-to-real RL for loco-manipulation typically involves complicated reward designs that are fine-tuned for specific environments and tasks [361, 362], see Sec. VII-A. To enable loco-manipulation tasks, these methods define the contact sequence either implicitly using reference trajectory [361, 344] or explicitly via reward design [281], thus increasing the success rate of sim-to-real transfer. To address the uncertainty of the mass and other properties of the manipulated object, most RL approaches rely on domain randomization to provide robustness to object parameters [42].
@@
A hierarchical RL structure that manages distinct skills is a viable strategy for achieving various loco-manipulation tasks such as falling recovery and ball kicking [282]. However, hierarchical RL has scalability issues, see Sec. VII-E. On the other hand, IL methods, particularly those employing RL with motion imitation, have enabled humanoid loco-manipulation via teleoperation, as outlined in Sec. VII-B. Although teleoperation is not autonomous, it is a crucial intermediate step to collect humanoid data. This has led to promising developments in autonomous loco-manipulation skills [267, 268], with the potential for further expansion to more diverse loco-manipulation tasks.
Key Takeaways: Although learning-based methods for humanoid loco-manipulation are less developed than model-based methods, their significance should not be overlooked. Learning-based methods are potentially more robust, as they can adapt to unstructured scenarios that model-based methods struggle to explicitly address, such as recovering from a fall in an arbitrary configuration [363]. Moreover, learning-based methods can find emergent behaviors that are challenging for model-based methods [291]. The recent success of quadrupedal loco-manipulation [86, 81, 85] shows significant promise for humanoids. However, transferring these algorithms from quadruped to humanoid is challenging because of the more complex dynamics of humanoids, which require enhanced safety measures and precise balance control.
VII-G Challenge in Lack of Benchmarks for Loco-manipulation
The development of humanoid loco-manipulation skills is still in its infancy compared to more established tasks such as humanoid locomotion [364] and tabletop manipulation [365]. To advance this emerging field, there is a pressing need for large-scale, systematic benchmarks specifically designed for humanoid loco-manipulation, encompassing both simulation and real-world environments. Benchmarks that define loco-manipulation tasks and evaluation metrics can significantly accelerate research in this field.
For benchmarks in simulation, HumanoidBench [286] and Mimicking-Bench [366] offer a list of loco-manipulation tasks, such as sitting on a piece of furniture and wiping a window. SMPLOlympics [367] presents a collection of sports as tasks for simulated humanoids. These benchmarks provide metrics for evaluating the performance of humanoid algorithms and help verify the reproducibility within MuJoCo [214] and Isaac Gym [215]. However, these tasks represent only a subset of human skills, and the evaluation metrics remain highly task-specific. Expanding the benchmarks to include a broader range of real-world tasks, along with more standardized metrics, would enhance their utility and practicality. Moreover, the sim-to-real gap continues to limit the practical value and reliability of these simulation benchmarks: it remains uncertain how well performance in simulation translates to real-world deployment, although promising results have been demonstrated in zero-shot sim-to-real transfer.
For hardware benchmarks, the development of affordable and capable humanoid robots as standardized platforms for hardware evaluation can significantly accelerate the research. Research efforts on open-source humanoid hardware and software, such as Hector [2], the MIT Humanoid [179], the Berkeley Humanoid [368], and the Duke Humanoid [46] represent valuable contributions. Additionally, rapid prototyping of diverse humanoid robots facilitates cross-embodiment generalization, while mass production of standardized robots streamlines the deployment of benchmarks, such as through the Robotic Grand Factory initiatives [369]. Complementing these efforts, the construction of remote-accessible real-world environments, as exemplified by the DARPA Robotics Challenge, Robotarium [370], and TriFinger [371], enables researchers to focus on advancing algorithms. Ideally, these physical environments share the same benchmark metrics as the simulators to facilitate sim-to-real transfer, while remaining highly customizable to support novel tasks.
VII-H Challenge in Data Scarcity
As discussed in Sec. VII-B, the four data sources present a trade-off between quality and availability. The lack of high-quality large-scale robot data becomes a bottleneck for robot skill learning. To solve the bottleneck, much effort has been put into data scaling. There is ongoing debate about whether scaling is the best approach to generalist humanoid robots.
We argue that generalization in loco-manipulation is achieved by including the motion of the manipulated objects, enhanced with a greater variety of sensing modalities instead of data quantity scaling. To develop truly versatile and adaptive humanoids, human data should also include that of the manipulated objects, cognitive actions (e.g., trust, compete, collaborate) paired with multimodal observations (e.g., whole-body haptic sensing, egocentric images), so that humanoids can learn the ‘how’ and ‘why’. However, instrumenting the environment and the manipulated objects with force and tactile sensing is highly complex and difficult to scale. Potential approaches include leveraging vision-language models to extract information from videos [373, 374, 375], employing video generation models [376] or physics-based simulation [214, 215] to generate synthetic data, and adopting hybrid strategies that effectively combine these synthetic data sources with limited real-world multimodal data [377, 378]. Recent work aims to bridge the gap between human animation and humanoid applications by collecting human multimodal observations [379, 380] and human kinetics [381]. Together with the rapid advancement in humanoid hardware, purposive learning with more informative human data will become the mainstream approach to achieve versatile and general-purpose humanoids.
@@ VIII Foundation Models for Humanoid Robots
Foundation Models (FMs) are large pre-trained models using Internet-scale data [382]. Recent progress in FMs such as Large Language Models (LLMs) and Vision-Language Models (VLMs) has demonstrated groundbreaking capabilities in solving a wide range of downstream tasks (through in-context learning or fine-tuning), such as code generation and video understanding [383]. The common sense reasoning capabilities of FMs have inspired many explorations of their applications in robotics [11, 12]. Despite this growing interest, research on FMs specifically for humanoid robots remains sparse. In this section, we first provide an overview of FMs in the context of general robotics (e.g., mobile manipulation, instead of humanoids) and then explore their potential applications to humanoid robots.
Strategies to leverage FMs for robotics can generally be categorized into two, as shown in Fig. 17. The first strategy (Sec. VIII-A) elicits actionable knowledge from pre-trained LLMs/VLMs for robotic tasks, without additional model fine-tuning. The second strategy (Sec. VIII-B) collects abundant robotic data to fine-tune or co-train a Robot Foundation Model (RFM) that generalizes to control tasks with common sense reasoning capability [384, 385, 386, 387, 388, 389].
@@ VIII-A Applying LLMs/VLMs to Humanoid Robots
Applying LLMs/VLMs to humanoids is still a nascent field. Many studies have demonstrated the successful deployment of LLMs/VLMs across various robot embodiments, such as dexterous hands [390], manipulators [391], mobile manipulators [392], quadrupedal robots [393], and bipedal robots [394].
@@
VIII-A1 FMs for selecting low-level skills (task planning)
Such task planning capability of FMs has advanced the complexity of tasks that a humanoid can accomplish. For instance,  [396] and [397] use pre-trained LLMs to select skills and task goals for animated humanoid characters. OmniH2O [267] employs GPT-4 [383] to select autonomous skills such as greeting a human. HYPERmotion [398] applies an LLM to construct task graphs that enable a hybrid wheeled-leg robot to execute complex loco-manipulation tasks. However, using FMs with a fixed skill set limits skill versatility. In addition, for complex behaviors such as those in humanoid loco-manipulation tasks, it is essential to allow FMs to author the detailed motions of low-level skills, instead of selecting from the existing, relatively-abstract low-level skills.
VIII-A2 FMs for authoring low-level skills
Thus, many research efforts have focused on identifying the best bridge between FMs and low-level robot skills. For example, researchers have proposed to generate code [391, 399] and reward functions [400, 393, 394] as intermediate representations for bipedal and quadrupedal robots. Compared to selecting the existing skills, these intermediate representations provide additional flexibility in adjusting the generated motion. Furthermore, FMs can generate whole-body poses [401, 402] and whole-body contacts [403, 404] for humanoid robots. These techniques allow users to intuitively direct a robot’s behavior through expressive inputs such as natural language, images, or even gestures.
@@ VIII-B Building Humanoid Foundation Models
While most FMs are developed in the vision or language domain, building them in the robotics domain for embodied intelligence is a natural extension. Similar to LLM and VLM, Robot Foundation Models (RFMs) are large models trained from Internet-scale robotic datasets. These RFMs often process multimodal inputs (e.g., egocentric images, and natural language as task description) and enable interaction with the physical world through robot action outputs. Therefore, they are also called Vision-Language-Action (VLA) models. Leveraging the Internet-scale multimodal dataset, RFMs hold the promise of generalization across diverse tasks and provide a natural interface for human-robot interaction, both essential for real-world robot applications.
VIII-B1 Hierarchical vision-language-conditioned policy
@@
One popular approach to building RFMs is to create a multi-task sensorimotor control policy using a high-capacity model that can consume large amounts of robot data. The trained policy can perform a wide variety of low-level skills, even for multiple robot embodiments. To enable high-level reasoning/planning, this approach often employs a hierarchical framework combining a pre-trained LLM or VLM with the low-level control policy, initialized from scratch. During training, this setup aligns the semantic knowledge of the LLM and VLM with the physical behavior of the control policy, enabling cross-modal capabilities such as language-to-action. Transformer [405] is a common choice for such a low-level policy due to its scalability. To enable interactive user commands, the low-level policy is conditioned on language and/or image inputs, often tokenized with pre-trained text or image encoders from LLMs or VLMs. For example, RT-1 [384] and VIMA [386] both train a language-conditioned visuomotor policy with a large amount of manipulation data and have demonstrated the ability to perform a wide range of skills. However, successful implementations are limited to robots that have stable dynamics and a large amount of high-quality data collected with significant resources [385, 406, 407, 387].
Applying such a hierarchical framework to humanoid robots is an exciting new direction. Recent works have extended RFMs to the humanoid upper body [377, 389, 388]. For example, NVIDIA developed GR00T N1 [377], a general-purpose FM for humanoid robots, and Figure AI demonstrates Helix, a hierarchical VLA for dexterous manipulation and collaboration between two humanoid robots. However, building an RFM for humanoid loco-manipulation remains a challenging endeavor: the inherent instability in locomotion and high-dimensional action space in dexterous hands makes it difficult to collect high-quality data efficiently. Therefore, VLA for humanoid loco-manipulation is shown only in simulation. For example, HumanVLA [312] and SuperPADL [313] train a humanoid action policy aligned with the latent space of pre-trained VLMs, enabling skills based on image and language inputs.
VIII-B2 End-to-end Vision-Language-Action model